For the function f(x)=3x-5 and g(x)=x2+1, find g(f(c))
Answer by stanbon(75887)

For the function f(x)=3x-5 and g(x)=x2+1,
find g(f(c)) = g[3c-5] = (3c-5)^2+1 = 9c^2-30c+25+1 = 9c^2-30c+26
==================
Cheers,
Stan H.